OpenStudy (anonymous):

(please show steps) solve the inequality: 1/x <(or equal to) 1/x+1

OpenStudy (anonymous):

OK, well, how do you solve an equation with a variable in the denominator?

OpenStudy (anonymous):

muahah. yay. actual not answering.

OpenStudy (anonymous):

\[1/x \le 1/x+1\]times all terms by x \[1 \le 1+x\]solve for x \[0 \le x\]\[x \ge 0\]

OpenStudy (anonymous):

is that a 1/x +1 or a 1/(x+1) ?

OpenStudy (anonymous):

(x+1)

OpenStudy (anonymous):

1/x-1/(x+1)<0 taking LCM you get 1/(x(x+1))<0 , This means -1<=x<=0

OpenStudy (anonymous):

Remember never cancel x on both sides of the inequality even if you get it common, take it the other side and factorize. (If x is negative the inequality is reversed on cancelling x n both sides)